Pringle Corporation acquired an 80% interest in Chip Corporation for $300,000 on January 1, 2012 when Chip's stockholders' equity consisted of $200,000 capital stock and $25,000 retained earnings. The excess cost over book value acquired was allocated to equipment that was undervalued by $50,000, inventory that was overvalued by $25,000 and to goodwill. The inventory was sold in 2012 and the equipment had a 5-year remaining useful life.

  • Chip regularly sells inventory to Pringle at 150% of cost. Intercompany sales were $120,000 in 2012 and $90,000 in 2013. Pringle's inventory included $30,000 of this merchandise at 12/31/12 and $45,000 of this merchandise at 12/31/13.

  • Pringle has $10,000 in accounts payable due to Chip.

    Required:

    Prepare the consolidation workpapers for Pringle Corporation using the process reviewed in class:

    1. Calculatetheunamortizeddifference/excess

    2. Calculate the goodwill or bargain purchase gain

    3. Calculate the unrealized profit for ending inventory for 2012 and 2013

    4. PreparethePurchasePriceAllocationandAmortizationSchedule

    5. Record all necessary elimination and adjusting journal entries

    6. Post the journal entries (as written in #5 above) to the Consolidation worksheet
Answer #1

1. Unrealised profit for ending inventory for 2012.

Sales price = $120,000

Sales Price = 150% of cost.

Hence Cost of sales = $120,000/150% = $80,000

Therefore, Profit for Chip on Sales to Pringle = 120,000-80,000 = 40,000

Therefore profit margin = 40,000/120,000 * 100 = 33.33% of Sales

Inventory on 31.12.2012 = 30,000

Therefore Unrealized profit = 30,000 * 33.33% = 10,000

Inventory on 31.12.2013 = 45,000

Therefore unrealized profit = 45,000 * 33.33% = 15,000

2. Calculation of Goodwill or Bargain Purchase Price :

Book Value of Chip Limited:

Capital Stock = 200,000.

Add: Retained Earnings = 25,000.

Add: Undervalued Stock = 50,000.

Less: Overvalued Inventory = 25,000.

Total = 200,000 + 25,000 + 50,000 - 25,000 = 250,000.

Therefore 100% Book Value = 250,000.

Accordingly 80% Book Value = 250,000 * 80% = 200,000.

Purchase Price = 300,000.

Therefore Goodwill = 300,000 - 200,000 = $ 100,000.

3. Journal Entries for elimination :

a) Accounts payable A/c. Dr. 10,000.

Accounts Receivable A/c Cr. 10,000.

Entry Date 31/12/12

b) Stock Reserve A/c Dr. 10,000

Inventory A/c Cr. 10,000

Entry Date 01/01/13

c) Inventory A/c Dr. 10,000

Stock Reserve A/c Cr. 10,000

Entry Date - 31/12/13

d) Stock Reserve A/c Dr. 15,000

Inventory A/c Cr. 15,000